OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of √128 feet and a width of √200 feet. Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form?

OpenStudy (anonymous):

i can help you break up sqrt 200 and sqrt of 128 into smaller numbers so \[\sqrt{200}\] becomes \[\sqrt{10 x 10 x 2}\] and \[\sqrt{128}\] b\[\sqrt{8x8x2}\]ecomes

OpenStudy (anonymous):

which equals 10\[\sqrt{2}\] and 8\[\sqrt{?}\]

OpenStudy (anonymous):

sorry, 10 sqrt 2 and 8 sqrt 2

OpenStudy (anonymous):

multiply those together and you get 80 sqrt 2

OpenStudy (anonymous):

and tahts the perimeter